OpenStudy (anonymous):

5. Multiply. x^2+8x+15/x-4 x^2-16/2x+6

OpenStudy (anonymous):

\[\frac{\left(x^2-16\right) \left(x^2+8 x+15\right)}{(x-4) (2 x+6)}=\frac{1}{2} (x+4) (x+5) \]